OpenStudy (anonymous):

how do you write f(x)=|2x+1|+1 as a piecewise function?

OpenStudy (anonymous):

oh i get to practice my piecewise in latex!

OpenStudy (anonymous):

ok so \[ |x| = \left\{\begin{array}{rcc} x & \text{if} & x \geq 0 \\ - x & \text{if} & x < 0 \end{array} \right. \]

OpenStudy (anonymous):

therefore \[ |2x+1| = \left\{\begin{array}{rcc} 2x +1& \text{if} & 2x+1 \geq 0 \\ - 2x-1 & \text{if} & 2x+1 < 0 \end{array} \right. \]

OpenStudy (anonymous):

but of course you want this in terms of x so you write \[ |2x+1| = \left\{\begin{array}{rcc} 2x +1& \text{if} & x \geq -\frac{1}{2} \\ - 2x-1 & \text{if} & x <- \frac{1}{2} \end{array} \right. \]

OpenStudy (anonymous):

and so \[ |2x+1| +1= \left\{\begin{array}{rcc} 2x +2& \text{if} & x \geq -\frac{1}{2} \\ - 2x & \text{if} & x <- \frac{1}{2} \end{array} \right. \]

OpenStudy (anonymous):

by adding one

OpenStudy (anonymous):

wow thanks!!! so if y=|4x+1|+2x-3 then the pieccewise function would be...

OpenStudy (anonymous):

|y|= 6x-2 if x\[\ge\] -1/4 -2x-4 if x\[>\]m1/4